# Deep Compression Autoencoder for Efficient High-Resolution Diffusion Models

### v5

@@ 4.1 Setups
For image generation experiments, we apply autoencoders to diffusion transformer models including DiT (Peebles & Xie, 2023) and UViT (Bao et al., 2023). We follow the same training settings as the original papers. Additionally, we build USiT by combining UViT (Bao et al., 2023) with the SiT sampler (Ma et al., 2024a). We consider three settings with different resolutions, including ImageNet (Deng et al., 2009) for generation, FFHQ (Karras et al., 2019) and MJHQ (Li et al., 2024a) for generation, and MapillaryVistas (Neuhold et al., 2017) for generation.
| Diffusion | Patch | Throughput (image/s) | Latency | Memory | FID | |||
| Model | Autoencoder | Size | Training | Inference | (ms) | (GB) | w/o CFG | w/ CFG |
| Flux-VAE-f8 [18] | 2 | 54 | 416 | 31.7 | 56.3 | 27.35 | 8.72 | |
| Asym-VAE-f8 [54] | 2 | 54 | 424 | 31.7 | 56.2 | 11.55 | 2.95 | |
| SD-VAE-f8 [37] | 2 | 54 | 424 | 31.7 | 56.2 | 12.03 | 3.04 | |
| DC-AE-f32 | 1 | 241 | 2016 | 7.8 | 20.9 | 9.56 | 2.84 | |
| DiT-XL [35] | DC-AE-f32‡ | 1 | 241 | 2016 | 7.8 | 20.9 | 6.88 | 2.41 |
| Flux-VAE-f8 [18] | 2 | 55 | 349 | 30.4 | 54.2 | 30.91 | 12.63 | |
| Asym-VAE-f8 [54] | 2 | 55 | 351 | 30.3 | 54.1 | 11.36 | 3.51 | |
| SD-VAE-f8 [37] | 2 | 55 | 351 | 30.3 | 54.1 | 11.04 | 3.55 | |
| DC-AE-f32 | 1 | 247 | 1622 | 8.2 | 18.6 | 9.83 | 2.53 | |
| DC-AE-f64 | 1 | 984 | 6706 | 3.5 | 10.6 | 13.96 | 3.01 | |
| UViT-H [1] | DC-AE-f64† | 1 | 984 | 6706 | 3.5 | 10.6 | 12.26 | 2.66 |
| Asym-VAE-f8 [54] | 2 | 27 | 157 | 74.8 | OOM | 9.87 | 3.62 | |
| SD-VAE-f8 [37] | 2 | 27 | 157 | 74.8 | OOM | 9.73 | 3.57 | |
| DC-AE-f32 | 1 | 112 | 665 | 19.7 | 42.0 | 8.13 | 2.30 | |
| DC-AE-f64 | 1 | 450 | 2733 | 8.6 | 30.2 | 7.78 | 2.47 | |
| UViT-2B [1] | DC-AE-f64† | 1 | 450 | 2733 | 8.6 | 30.2 | 6.50 | 2.25 |
| MAGVIT-v2 [48] | - | - | - | - | - | - | 3.07 | 1.91 |
| EDM2-XXL [15] | SD-VAE-f8 [37] | - | - | - | - | - | 1.91 | 1.81 |
| MAR-L [22] | - | - | - | - | - | - | 2.74 | 1.73 |
| USiT-H | DC-AE-f32 | 1 | 247 | 1622 | 8.2 | 18.6 | 3.80 | 1.89 |
| USiT-2B | DC-AE-f32 | 1 | 112 | 665 | 19.7 | 42.0 | 2.90 | 1.72 |
| Diffusion | Patch | Throughput (image/s) | Latency | Memory | MJHQ 512512 | |||
| Model | Autoencoder | Size | Training | Inference | (ms) | (GB) | FID | CLIP Score |
| SD-VAE-f8 [37] | 2 | 43 | 312 | 37.1 | 60.45 | 6.3 | 26.36 | |
| PIXART- [6] | DC-AE-f32 | 1 | 173 | 1251 | 10.4 | 23.77 | 6.1 | 26.41 |
@@
We profile the training and inference throughput on the H100 GPU with PyTorch and TensorRT respectively. The latency is measured on the 3090 GPU with batch size 2. The training memory is profiled using PyTorch, assuming a batch size of 256. We use fp16 for all cases. For simplicity, we assume the number of sampling steps is 1.
@@ Acknowledgements
Appendix A Additional Samples
Appendix B Diffusion Model Architecture Details
@@
Appendix C High-Resolution Image Generation Results
| FFHQ 10241024 (Unconditional) & MJHQ 10241024 (Class-Conditional) | |||||||||
| Diffusion | Patch | Throughput (image/s) | Latency | Memory | FFHQ FID | MJHQ FID | |||
| Model | Autoencoder | Size | Training | Inference | (ms) | (GB) | w/o CFG | w/o CFG | w/ CFG |
| SD3-VAE-f8 [9] | 2 | 83 | 814 | 14.2 | 41.4 | 46.28 | 109.43 | 103.02 | |
| Flux-VAE-f8 [18] | 2 | 83 | 814 | 14.2 | 41.4 | 59.15 | 143.16 | 139.06 | |
| SDXL-VAE-f8 [36] | 2 | 84 | 833 | 14.1 | 41.2 | 16.82 | 49.00 | 39.21 | |
| Asym-VAE-f8 [54] | 2 | 84 | 833 | 14.1 | 41.2 | 17.12 | 48.25 | 38.36 | |
| 2 | 84 | 833 | 14.1 | 41.2 | 16.98 | 48.05 | 38.19 | ||
| SD-VAE-f8 [37] | 4 | 470 | 5566 | 2.5 | 10.7 | 23.81 | 60.94 | 51.29 | |
| DC-AE-f32 | 1 | 475 | 5575 | 2.5 | 10.7 | 13.65 | 34.35 | 27.20 | |
| DC-AE-f32‡ | 1 | 475 | 5575 | 2.5 | 10.7 | 11.39 | 28.36 | 21.89 | |
| DiT-S [35] | DC-AE-f64 | 1 | 2085 | 25259 | 1.0 | 3.1 | 26.88 | 61.30 | 53.38 |
| MapillaryVistas 20482048 (Unconditional) | |||||||||
| Diffusion | Patch | Throughput (image/s) | Latency | Memory | MapillaryVistas FID | ||||
| Model | Autoencoder | Size | Training | Inference | (ms) | (GB) | w/o CFG | ||
| SD-VAE-f8 [37] | 4 | 84 | 810 | 14.3 | 41.4 | 69.50 | |||
| DiT-S [35] | DC-AE-f64 | 1 | 459 | 5435 | 2.6 | 11.0 | 59.55 | ||
@@ Appendix C High-Resolution Image Generation Results
Appendix D Image Generation Results with Other Evaluation Metrics
@@
| Diffusion | Patch | Inference | FID | Inception Score | Precision | Recall | CMMD | ||||||
| Model | Autoencoder | Size | Throughput | w/o CFG | w/ CFG | w/o CFG | w/ CFG | w/o CFG | w/ CFG | w/o CFG | w/ CFG | w/o CFG | w/ CFG |
| SD3-VAE-f8 [9] | 2 | 2984 | 164.34 | 143.82 | 6.07 | 7.53 | 0.06 | 0.09 | 0.31 | 0.39 | 3.13 | 2.94 | |
| Flux-VAE-f8 [18] | 2 | 2984 | 106.07 | 84.73 | 13.39 | 17.71 | 0.28 | 0.37 | 0.39 | 0.42 | 1.90 | 1.67 | |
| SDXL-VAE-f8 [36] | 2 | 2991 | 51.03 | 26.38 | 27.58 | 56.72 | 0.57 | 0.74 | 0.58 | 0.50 | 1.35 | 1.05 | |
| Asym-VAE-f8 [54] | 2 | 2991 | 52.68 | 25.14 | 30.22 | 65.27 | 0.58 | 0.74 | 0.62 | 0.51 | 1.09 | 0.80 | |
| SD-VAE-f8 [37] | 2 | 2991 | 51.96 | 24.57 | 30.37 | 65.73 | 0.57 | 0.74 | 0.64 | 0.52 | 1.23 | 0.91 | |
| SD-VAE-f16 [37] | 2 | 12881 | 76.86 | 44.22 | 21.38 | 43.35 | 0.43 | 0.62 | 0.60 | 0.55 | 1.83 | 1.46 | |
| SD-VAE-f32 [37] | 1 | 12883 | 70.23 | 38.63 | 23.07 | 47.72 | 0.46 | 0.64 | 0.58 | 0.56 | 1.71 | 1.36 | |
| DC-AE-f32 | 1 | 12850 | 46.12 | 18.08 | 34.82 | 84.73 | 0.59 | 0.76 | 0.66 | 0.56 | 1.00 | 0.70 | |
| DC-AE-f64 | 1 | 53774 | 67.30 | 35.96 | 24.55 | 52.86 | 0.44 | 0.64 | 0.60 | 0.56 | 1.44 | 1.14 | |
| UViT-S [1] | DC-AE-f64† | 1 | 53774 | 61.84 | 30.63 | 27.28 | 61.76 | 0.47 | 0.67 | 0.63 | 0.56 | 1.35 | 1.04 |
| Flux-VAE-f8 [18] | 2 | 416 | 27.35 | 8.72 | 53.09 | 130.20 | 0.68 | 0.83 | 0.61 | 0.48 | 0.54 | 0.30 | |
| Asym-VAE-f8 [54] | 2 | 424 | 11.39 | 2.97 | 108.70 | 241.10 | 0.75 | 0.83 | 0.65 | 0.53 | 0.37 | 0.20 | |
| SD-VAE-f8 [37] | 2 | 424 | 12.03 | 3.04 | 105.25 | 240.82 | 0.75 | 0.84 | 0.64 | 0.54 | 0.43 | 0.25 | |
| DC-AE-f32 | 1 | 2016 | 9.56 | 2.84 | 117.49 | 226.98 | 0.75 | 0.82 | 0.64 | 0.55 | 0.34 | 0.22 | |
| DiT-XL [35] | DC-AE-f32‡ | 1 | 2016 | 6.88 | 2.41 | 141.07 | 263.56 | 0.76 | 0.82 | 0.63 | 0.56 | 0.29 | 0.18 |
| Flux-VAE-f8 [18] | 2 | 349 | 30.91 | 12.63 | 56.72 | 127.93 | 0.64 | 0.76 | 0.59 | 0.49 | 0.50 | 0.31 | |
| Asym-VAE-f8 [54] | 2 | 351 | 11.36 | 3.51 | 124.24 | 249.21 | 0.75 | 0.82 | 0.61 | 0.53 | 0.32 | 0.20 | |
| SD-VAE-f8 [37] | 2 | 351 | 11.04 | 3.55 | 125.08 | 250.66 | 0.75 | 0.82 | 0.61 | 0.53 | 0.39 | 0.26 | |
| DC-AE-f32 | 1 | 1622 | 9.83 | 2.53 | 121.91 | 255.07 | 0.76 | 0.83 | 0.65 | 0.54 | 0.34 | 0.20 | |
| DC-AE-f64 | 1 | 6706 | 13.96 | 3.01 | 99.20 | 229.16 | 0.73 | 0.83 | 0.64 | 0.53 | 0.50 | 0.31 | |
| UViT-H [1] | DC-AE-f64† | 1 | 6706 | 12.26 | 2.66 | 109.20 | 239.82 | 0.73 | 0.82 | 0.67 | 0.57 | 0.43 | 0.27 |
| Flux-VAE-f8 [18] | 2 | 155 | 25.03 | 10.12 | 74.04 | 161.29 | 0.67 | 0.78 | 0.58 | 0.51 | 0.38 | 0.24 | |
| Asym-VAE-f8 [54] | 2 | 157 | 9.87 | 3.62 | 131.95 | 258.63 | 0.76 | 0.83 | 0.59 | 0.52 | 0.30 | 0.19 | |
| SD-VAE-f8 [37] | 2 | 157 | 9.73 | 3.57 | 132.86 | 260.50 | 0.76 | 0.83 | 0.59 | 0.52 | 0.37 | 0.24 | |
| DC-AE-f32 | 1 | 665 | 8.13 | 2.30 | 135.44 | 272.73 | 0.76 | 0.82 | 0.66 | 0.56 | 0.30 | 0.17 | |
| DC-AE-f64 | 1 | 2733 | 7.78 | 2.47 | 138.11 | 280.49 | 0.77 | 0.84 | 0.63 | 0.54 | 0.35 | 0.22 | |
| UViT-2B [1] | DC-AE-f64† | 1 | 2733 | 6.50 | 2.25 | 152.35 | 293.45 | 0.77 | 0.83 | 0.65 | 0.56 | 0.31 | 0.19 |
| USiT-H | DC-AE-f32 | 1 | 1622 | 3.80 | 1.89 | 174.58 | 252.35 | 0.78 | 0.82 | 0.64 | 0.60 | 0.24 | 0.18 |
| USiT-2B | DC-AE-f32 | 1 | 665 | 2.90 | 1.72 | 187.68 | 248.10 | 0.79 | 0.82 | 0.63 | 0.61 | 0.21 | 0.17 |

### v6

@@ 4.1 Setups
For image generation experiments, we apply autoencoders to diffusion transformer models including DiT (Peebles & Xie, 2023) and UViT (Bao et al., 2023). We follow the same training settings as the original papers. Additionally, we build USiT by combining UViT (Bao et al., 2023) with the SiT sampler (Ma et al., 2024a). The SiT and USiT models are trained for 500k iterations with batch size 1024. We consider three settings with different resolutions, including ImageNet (Deng et al., 2009) for generation, FFHQ (Karras et al., 2019) and MJHQ (Li et al., 2024a) for generation, and MapillaryVistas (Neuhold et al., 2017) for generation.
Diffusion | Patch | Throughput (image/s) | Latency | Memory | FID | ||||
Model | Autoencoder | Size | NFE | Training | Inference | (ms) | (GB) | w/o CFG | w/ CFG |
Flux-VAE-f8 [20] | 2 | 250 | 54 | 0.83 | 7915 | 56.3 | 27.35 | 8.72 | |
Asym-VAE-f8 [58] | 2 | 250 | 54 | 0.85 | 7686 | 56.2 | 11.39 | 2.97 | |
SD-VAE-f8 [40] | 2 | 250 | 54 | 0.85 | 7686 | 56.2 | 12.03 | 3.04 | |
DC-AE-f32 | 1 | 250 | 241 | 4.03 | 1958 | 20.9 | 9.56 | 2.84 | |
DiT-XL [38] | DC-AE-f32‡ | 1 | 250 | 241 | 4.03 | 1958 | 20.9 | 6.88 | 2.41 |
Flux-VAE-f8 [20] | 2 | 30 | 55 | 5.82 | 913 | 54.2 | 30.91 | 12.63 | |
Asym-VAE-f8 [58] | 2 | 30 | 55 | 5.85 | 914 | 54.1 | 11.36 | 3.51 | |
SD-VAE-f8 [40] | 2 | 30 | 55 | 5.85 | 914 | 54.1 | 11.04 | 3.55 | |
DC-AE-f32 | 1 | 30 | 247 | 27.03 | 246 | 18.6 | 9.83 | 2.53 | |
DC-AE-f64 | 1 | 30 | 984 | 111.77 | 104 | 10.6 | 13.96 | 3.01 | |
UViT-H [1] | DC-AE-f64† | 1 | 30 | 984 | 111.77 | 105 | 10.6 | 12.26 | 2.66 |
Asym-VAE-f8 [58] | 2 | 30 | 27 | 2.62 | 2243 | OOM | 9.87 | 3.62 | |
SD-VAE-f8 [40] | 2 | 30 | 27 | 2.62 | 2243 | OOM | 9.73 | 3.57 | |
DC-AE-f32 | 1 | 30 | 112 | 11.08 | 590 | 42.0 | 8.13 | 2.30 | |
DC-AE-f64 | 1 | 30 | 450 | 45.55 | 258 | 30.2 | 7.78 | 2.47 | |
UViT-2B [1] | DC-AE-f64† | 1 | 30 | 450 | 45.55 | 258 | 30.2 | 6.50 | 2.25 |
MAGVIT-v2 [51] | - | - | - | - | - | - | - | 3.07 | 1.91 |
EDM2-XXL [17] | - | - | - | - | - | - | - | 1.91 | 1.81 |
MAR-L [24] | - | - | - | - | - | - | - | 2.74 | 1.73 |
SiT-XL [33] | DC-AE-f32 | 1 | - | 241 | - | - | 20.9 | 7.47 | 2.41 |
USiT-H | DC-AE-f32 | 1 | - | 247 | - | - | 18.6 | 3.80 | 1.89 |
USiT-2B | DC-AE-f32 | 1 | - | 112 | - | - | 42.0 | 2.90 | 1.72 |
Diffusion | Patch | Throughput (image/s) | Latency | Memory | MJHQ 512512 | ||||
Model | Autoencoder | Size | NFE | Training | Inference | (ms) | (GB) | FID | CLIP Score |
SD-VAE-f8 [40] | 2 | 20 | 43 | 7.81 | 742 | 60.45 | 6.3 | 26.36 | |
PIXART- [6] | DC-AE-f32 | 1 | 20 | 173 | 31.27 | 209 | 23.77 | 6.1 | 26.41 |
@@
We profile the training and inference throughput on the H100 GPU with PyTorch and TensorRT respectively. The latency is measured on the 3090 GPU with batch size 2. The training memory is profiled using PyTorch, assuming a batch size of 256. We use fp16 for all cases.
@@ Acknowledgements
Appendix A DC-AE Architecture and Training Details
We present the detailed architecture of DC-AE encoder and decoder stages in Figure 10 to complement Figure 4 (b).
We use the AdamW optimizer (Loshchilov, 2017) for all training phases.
In phase 1 (low-resolution full training), we use a constant learning rate of 6.4e-5 with a weight decay of 0.1, and AdamW betas of (0.9, 0.999). We use L1 loss and LPIPS loss (Zhang et al., 2018).
In phase 2 (high-resolution latent adaptation), we use a constant learning rate of 1.6e-5, a weight decay of 0.001, and AdamW betas of (0.9, 0.999). We use the same loss as phase 1.
Appendix B Ablation Study on Training Different Numbers of Layers
Figure 11 presents the ablation study on training different numbers of layers in phase 2 (high-resolution latent adaptation) and phase 3 (low-resolution local refinement).
Appendix C Additional Image Reconstruction Results
Table 5 reports the reconstruction results under the low spatial-compression ratio setting. DC-AE delivers slightly better results than SD-VAE under this setting.
f8c4 | 32324 | SD-VAE [40] | 0.63 | 24.99 | 0.71 | 0.063 |
DC-AE | 0.46 | 25.46 | 0.73 | 0.057 |
Appendix D Latent Scaling and Shifting Factors
Following the common practice (Rombach et al., 2022; Peebles & Xie, 2023; Bao et al., 2023; Esser et al., 2024; Labs, 2024; Chen et al., 2024b; a), we normalize the latent space of our autoencoders to apply to latent diffusion models. Given a dataset, we compute the root mean square of the latent features and use its multiplicative inverse as the scaling factor for our autoencoders. We do not use the shifting factor for our autoencoders.
Appendix E Diffusion Model Architecture Details
@@
Appendix F Diffusion Sampling Hyperparameters
For the DiT models, we use the DDPM (Ho et al., 2020) sampler from the DiT (Peebles & Xie, 2023) codebase with 250 sampling steps and a guidance scale of 1.3.
For the UViT models, we use the DPMSolver (Lu et al., 2022a) sampler with 30 sampling steps and a guidance scale of 1.5.
Appendix G High-Resolution Image Generation Results
FFHQ 10241024 (Unconditional) & MJHQ 10241024 (Class-Conditional) | ||||||||||
Diffusion | Patch | Throughput (image/s) | Latency | Memory | FFHQ FID | MJHQ FID | ||||
Model | Autoencoder | Size | NFE | Training | Inference | (ms) | (GB) | w/o CFG | w/o CFG | w/ CFG |
SD3-VAE-f8 [9] | 2 | 250 | 83 | 1.63 | 3554 | 41.4 | 46.28 | 109.43 | 103.02 | |
Flux-VAE-f8 [20] | 2 | 250 | 83 | 1.63 | 3554 | 41.4 | 59.15 | 143.16 | 139.06 | |
SDXL-VAE-f8 [39] | 2 | 250 | 84 | 1.67 | 3530 | 41.2 | 16.82 | 49.00 | 39.21 | |
Asym-VAE-f8 [58] | 2 | 250 | 84 | 1.67 | 3530 | 41.2 | 17.10 | 48.30 | 38.35 | |
2 | 250 | 84 | 1.67 | 3530 | 41.2 | 16.98 | 48.05 | 38.19 | ||
SD-VAE-f8 [40] | 4 | 250 | 470 | 11.13 | 632 | 10.7 | 23.81 | 60.94 | 51.29 | |
DC-AE-f32 | 1 | 250 | 475 | 11.15 | 634 | 10.7 | 13.65 | 34.35 | 27.20 | |
DC-AE-f32‡ | 1 | 250 | 475 | 11.15 | 634 | 10.7 | 11.39 | 28.36 | 21.89 | |
DiT-S [38] | DC-AE-f64 | 1 | 250 | 2085 | 50.26 | 230 | 3.1 | 26.88 | 61.30 | 53.38 |
MapillaryVistas 20482048 (Unconditional) | ||||||||||
Diffusion | Patch | Throughput (image/s) | Latency | Memory | MapillaryVistas FID | |||||
Model | Autoencoder | Size | NFE | Training | Inference | (ms) | (GB) | w/o CFG | ||
SD-VAE-f8 [40] | 4 | 250 | 84 | 1.64 | 3561 | 41.4 | 69.50 | |||
DiT-S [38] | DC-AE-f64 | 1 | 250 | 459 | 10.91 | 639 | 11.0 | 59.55 |
@@ Appendix G High-Resolution Image Generation Results
Appendix H Image Generation Results with Other Evaluation Metrics
@@
Diffusion | Patch | Inference | FID | Inception Score | Precision | Recall | CMMD | |||||||
Model | Autoencoder | Size | NFE | Throughput | w/o CFG | w/ CFG | w/o CFG | w/ CFG | w/o CFG | w/ CFG | w/o CFG | w/ CFG | w/o CFG | w/ CFG |
SD3-VAE-f8 [9] | 2 | 30 | 49.73 | 164.34 | 143.82 | 6.07 | 7.53 | 0.06 | 0.09 | 0.31 | 0.39 | 3.13 | 2.94 | |
Flux-VAE-f8 [20] | 2 | 30 | 49.73 | 106.07 | 84.73 | 13.39 | 17.71 | 0.28 | 0.37 | 0.39 | 0.42 | 1.90 | 1.67 | |
SDXL-VAE-f8 [39] | 2 | 30 | 49.85 | 51.03 | 26.38 | 27.58 | 56.72 | 0.57 | 0.74 | 0.58 | 0.50 | 1.35 | 1.05 | |
Asym-VAE-f8 [58] | 2 | 30 | 49.85 | 52.68 | 25.14 | 30.22 | 65.27 | 0.58 | 0.74 | 0.62 | 0.51 | 1.09 | 0.80 | |
SD-VAE-f8 [40] | 2 | 30 | 49.85 | 51.96 | 24.57 | 30.37 | 65.73 | 0.57 | 0.74 | 0.64 | 0.52 | 1.23 | 0.91 | |
SD-VAE-f16 [40] | 2 | 30 | 214.68 | 76.86 | 44.22 | 21.38 | 43.35 | 0.43 | 0.62 | 0.60 | 0.55 | 1.83 | 1.46 | |
SD-VAE-f32 [40] | 1 | 30 | 214.72 | 70.23 | 38.63 | 23.07 | 47.72 | 0.46 | 0.64 | 0.58 | 0.56 | 1.71 | 1.36 | |
DC-AE-f32 | 1 | 30 | 214.17 | 46.12 | 18.08 | 34.82 | 84.73 | 0.59 | 0.76 | 0.66 | 0.56 | 1.00 | 0.70 | |
DC-AE-f64 | 1 | 30 | 896.23 | 67.30 | 35.96 | 24.55 | 52.86 | 0.44 | 0.64 | 0.60 | 0.56 | 1.44 | 1.14 | |
UViT-S [1] | DC-AE-f64† | 1 | 30 | 896.23 | 61.84 | 30.63 | 27.28 | 61.76 | 0.47 | 0.67 | 0.63 | 0.56 | 1.35 | 1.04 |
Flux-VAE-f8 [20] | 2 | 250 | 0.83 | 27.35 | 8.72 | 53.09 | 130.20 | 0.68 | 0.83 | 0.61 | 0.48 | 0.54 | 0.30 | |
Asym-VAE-f8 [58] | 2 | 250 | 0.85 | 11.39 | 2.97 | 108.70 | 241.10 | 0.75 | 0.83 | 0.65 | 0.53 | 0.37 | 0.20 | |
SD-VAE-f8 [40] | 2 | 250 | 0.85 | 12.03 | 3.04 | 105.25 | 240.82 | 0.75 | 0.84 | 0.64 | 0.54 | 0.43 | 0.25 | |
DC-AE-f32 | 1 | 250 | 4.03 | 9.56 | 2.84 | 117.49 | 226.98 | 0.75 | 0.82 | 0.64 | 0.55 | 0.34 | 0.22 | |
DiT-XL [38] | DC-AE-f32‡ | 1 | 250 | 4.03 | 6.88 | 2.41 | 141.07 | 263.56 | 0.76 | 0.82 | 0.63 | 0.56 | 0.29 | 0.18 |
Flux-VAE-f8 [20] | 2 | 30 | 5.82 | 30.91 | 12.63 | 56.72 | 127.93 | 0.64 | 0.76 | 0.59 | 0.49 | 0.50 | 0.31 | |
Asym-VAE-f8 [58] | 2 | 30 | 5.85 | 11.36 | 3.51 | 124.24 | 249.21 | 0.75 | 0.82 | 0.61 | 0.53 | 0.32 | 0.20 | |
SD-VAE-f8 [40] | 2 | 30 | 5.85 | 11.04 | 3.55 | 125.08 | 250.66 | 0.75 | 0.82 | 0.61 | 0.53 | 0.39 | 0.26 | |
DC-AE-f32 | 1 | 30 | 27.03 | 9.83 | 2.53 | 121.91 | 255.07 | 0.76 | 0.83 | 0.65 | 0.54 | 0.34 | 0.20 | |
DC-AE-f64 | 1 | 30 | 111.77 | 13.96 | 3.01 | 99.20 | 229.16 | 0.73 | 0.83 | 0.64 | 0.53 | 0.50 | 0.31 | |
UViT-H [1] | DC-AE-f64† | 1 | 30 | 111.77 | 12.26 | 2.66 | 109.20 | 239.82 | 0.73 | 0.82 | 0.67 | 0.57 | 0.43 | 0.27 |
Flux-VAE-f8 [20] | 2 | 30 | 2.58 | 25.03 | 10.12 | 74.04 | 161.29 | 0.67 | 0.78 | 0.58 | 0.51 | 0.38 | 0.24 | |
Asym-VAE-f8 [58] | 2 | 30 | 2.62 | 9.87 | 3.62 | 131.95 | 258.63 | 0.76 | 0.83 | 0.59 | 0.52 | 0.30 | 0.19 | |
SD-VAE-f8 [40] | 2 | 30 | 2.62 | 9.73 | 3.57 | 132.86 | 260.50 | 0.76 | 0.83 | 0.59 | 0.52 | 0.37 | 0.24 | |
DC-AE-f32 | 1 | 30 | 11.08 | 8.13 | 2.30 | 135.44 | 272.73 | 0.76 | 0.82 | 0.66 | 0.56 | 0.30 | 0.17 | |
DC-AE-f64 | 1 | 30 | 45.55 | 7.78 | 2.47 | 138.11 | 280.49 | 0.77 | 0.84 | 0.63 | 0.54 | 0.35 | 0.22 | |
UViT-2B [1] | DC-AE-f64† | 1 | 30 | 45.55 | 6.50 | 2.25 | 152.35 | 293.45 | 0.77 | 0.83 | 0.65 | 0.56 | 0.31 | 0.19 |
USiT-H | DC-AE-f32 | 1 | - | - | 3.80 | 1.89 | 174.58 | 252.35 | 0.78 | 0.82 | 0.64 | 0.60 | 0.24 | 0.18 |
USiT-2B | DC-AE-f32 | 1 | - | - | 2.90 | 1.72 | 187.68 | 248.10 | 0.79 | 0.82 | 0.63 | 0.61 | 0.21 | 0.17 |
@@
Appendix I Additional Samples
In Figure 13 and 14, we provide additional image reconstruction samples produced by SD-VAE and DC-AE. Reconstructed images by DC-AE demonstrate better visual qualities than SD-VAE’s reconstructed images, especially for the f64 and f128 autoencoders. Some samples are cropped for better visualization of details like human faces and small texts.